The following is a chart of 25 baseball players' salaries and statistics from 2016.

Player Name RBI's HR's AVG Salary (in millions)
Joe Mauer 49 11 0.261 23.000
Robinson Cano 103 39 0.298 24.050
Leonys Martin 47 15 0.245 4.150
Brandon Crawford 84 12 0.275 6.000
Colby Rasmus 54 15 0.206 15.800
Carlos Gonzalez 100 25 0.298 17.454
Matt Kemp 108 35 0.268 21.500
Prince Fielder 44 8 0.212 18.000
Mark Teixeira 44 15 0.204 23.125
Yoenis Cespedes 86 31 0.284 27.500
Chris Iannetta 24 7 0.210 4.550
Ryan Howard 59 25 0.196 25.000
Matt Wieters 66 17 0.243 15.800
Jayson Werth 70 21 0.244 21.571
Justin Smoak 34 14 0.217 3.900
Adrian Gonzalez 90 18 0.285 21.857
Coco Crisp 55 13 0.231 11.000
Ben Zobrist 76 18 0.272 10.500
J.D. Martinez 68 22 0.307 6.750
Aaron Hill 38 10 0.262 12.000
Adrian Beltre 104 32 0.300 18.000
David Ortiz 127 38 0.315 16.000
Chris Davis 84 38 0.221 21.119
Evan Gattis 72 32 0.251 3.300
Curtis Granderson 59 30 0.237 16.000



In order to have correlation with 95% significance, what is the critical r-value that we would like to have?  

(Round to three decimal places for all answers on this assignment.)

RBI vs. Salary

Complete a correlation analysis, using RBI's as the x-value and salary as the y-value.

Correlation coefficient:

Regression Equation: y=y=    

Do you have significant correlation? Select an answer Yes No

HR vs. Salary

Complete a correlation analysis, using HR's as the x-value and salary as the y-value.

Correlation coefficient:

Regression Equation: y=y=       

Do you have significant correlation? Select an answer Yes No   

AVG vs. Salary

Complete a correlation analysis, using AVG as the x-value and salary as the y-value.

Correlation coefficient:   

Regression Equation: y=y=       

Do you have significant correlation? Select an answer Yes No

Prediction

Based on your analysis, if you had to predict a player's salary, which method would be the best? Select an answer Regression equation with RBI's Regression equation with HR's Regression equation with AVG The average of the 25 salaries

Using that method, predict the salary for Ryan Braun. His stats were:

RBI: 91

HR: 31

AVG: 0.305

Based on your analysis, his predicted salary would be: $ million  

His actual salary was $20.000 million.
Answer #1

df = n-2 = 25-2 = 23

critical r for 95% confidence level =

0.413

RBI's HR's AVG Salary (in millions)
RBI's 1
HR's 0.783324 1
AVG 0.745874 0.444824 1
Salary (in millions) 0.369887 0.384432 0.05653 1

RBI vs. Salary

Complete a correlation analysis, using RBI's as the x-value and salary as the y-value.

Correlation coefficient: 0.369887

Regression Equation: y^ = 8.1429 + 0.1056 RBI

Do you have significant correlation? No

as r < critical value (0.413)

HR vs. Salary

Complete a correlation analysis, using HR's as the x-value and salary as the y-value.

Correlation coefficient: 0.783324

Regression Equation: y^ = 9.3843 + 0.2834 HR

Do you have significant correlation? yes

r > critical r

AVG vs. Salary

Complete a correlation analysis, using AVG as the x-value and salary as the y-value.

Correlation coefficient:    0.745874

Regression Equation: y^ = 12.5309 + 11.7711 AVG

Do you have significant correlation? yes

r > critical r

Prediction

Based on your analysis, if you had to predict a player's salary, which method would be the best?   Regression equation with HR's

Using that method, predict the salary for Ryan Braun. His stats were:

RBI: 91

HR: 31

AVG: 0.305

y^ = 9.3843 + 0.2834 *31 = 18.1697

Based on your analysis, his predicted salary would be: $ 18.1697 million  

His actual salary was $20.000 million.